### v3

@@ Abstract
When learning stable linear dynamical systems from data, three important properties are desirable: i) predictive accuracy, ii) provable stability, and iii) computational efficiency. Unconstrained minimization of reconstruction errors leads to high accuracy and efficiency but cannot guarantee stability. Existing methods to remedy this focus on enforcing stability while also ensuring accuracy, but do so only at the cost of increased computation. In this work, we investigate if a straightforward approach can simultaneously offer all three desiderata of learning stable linear systems. Specifically, we consider a post-hoc approach that manipulates the spectrum of the learned system matrix after it is learned in an unconstrained fashion. We call this approach spectrum clipping (SC) as it involves eigen decomposition and subsequent reconstruction of the system matrix after clipping all of its eigenvalues that are larger than one to one (without altering the eigenvectors). Through detailed experiments involving two different applications and publicly available benchmark datasets, we demonstrate that this simple technique can simultaneously learn highly accurate linear systems that are provably stable. Notably, we demonstrate that SC can achieve similar or better performance than strong baselines while being orders-of-magnitude faster. We also show that SC can be readily combined with Koopman operators to learn stable nonlinear dynamics, such as those underlying complex dexterous manipulation skills involving multi-fingered robotic hands. Further, we find that SC can learn stable robot policies even when the training data includes unsuccessful or truncated demonstrations. Our codes and dataset can be found at https://github.com/GT-STAR-Lab/spec_clip.
keywords:
Linear Dynamical Systems, Stability, System Identification.
@@ 1 Introduction
In spite of the growing dominance of deep neural networks in various fields, modeling and learning linear dynamical systems remain relevant due to the availability for analytical solutions, significantly lower computational burden, and ease of analyses and inspection. They continue to find use in diverse applications such as computer vision (Chan and Vasconcelos, 2005; Boots et al., 2007) and time series prediction (Liu et al., 2024; Takeishi et al., 2017). Even when the underlying dynamics are nonlinear, researchers are increasingly utilizing tools from linear systems theory by leveraging the Koopman Operator theory (Koopman, 1931; Williams et al., 2015) to approximately represent nonlinear systems as linear systems in a higher-dimensional lifted state space. These techniques have further increased the relevance and importance of learning linear systems in complex applications, such as differential drive robots (Abraham et al., 2017), serial-link manipulators (Abraham and Murphey, 2019), legged robots (Kim et al., 2024; Li et al., 2024), soft robotic manipulators (Bruder et al., 2021), and dexterous robotic manipulation (Han et al., 2023, 2024; Chen et al., 2024).
@@
When time-series data associated with system evolution is available, learning such linear systems can be considered as a form of self-supervised learning and can be solved efficiently by minimizing a regression objective (Ljung, ). However, such unconstrained optimization methods tend to overlook the important properties of the underlying linear systems, such as stability. This is particularly critical in linear system modeling, as a linear system comprises only three evolution modes: convergence, divergence, and oscillation (Robinson, 2012). Consequently, even if the underlying system dynamics are stable, the stability of learned system could be highly sensitive to the training data distribution, potentially leading to unstable solutions and catastrophic long-term predictions.
@@
A number of methods have been developed to enforce stability of learned linear systems and reduce sensitivity to training data (Lacy and Bernstein, 2002, 2003; Boots et al., 2007; Huang et al., 2016; Gillis et al., 2020; Mamakoukas et al., 2020, 2023). These methods take one of three main approaches: a) minimizing the regression objective under the stability constraints (Lacy and Bernstein, 2002, 2003), b) iterating between solving an unconstrained problem and optimizing its solution under the stability constraints (Boots et al., 2007; Huang et al., 2016), and c). iterating between characterizing the stable matrix and optimizing the characterizations with the regression objective (Gillis et al., 2020; Mamakoukas et al., 2020, 2023). Regardless of the underlying approach, these methods inevitably increase the computation burden, potentially counteracting one of the primary advantages of leveraging linear techniques. Furthermore, they exhibit poor scalability when applied to high-dimensional systems, which are necessary in both modeling inherently high-dimensional state spaces (e.g., images) and effectively approximating complex nonlinear systems using Koopman operators Koopman (1931) (e.g., robotic systems).
@@
In this work, we suggest and investigate a simple yet surprisingly effective method, Spectrum Clipping (SC), to learn the stable linear systems from data requiring negligible additional computations. This approach involves a straightforward implementation: 1) compute the system evolution matrix by solving the least-squares regression problem, 2) perform eigen decomposition of the system matrix and extract all eigenvalues, 3) clip eigenvalues with magnitudes larger than one to one, and 4) reconstruct the system evolution matrix using the clipped eigenvalues and the unchanged eigenvectors. SC enforces stability by construction but neither requires complex constrained optimization nor updating the matrix iteratively. As a result, SC is both efficient and scales well to high-dimensional systems. We are inspired to investigate this simple approach since strategies related to clipping a matrix’s spectrum have proven effective in other domains. For instance, directly controlling the singular values of linear layers seems to enhance training stability and reduce the model’s generalization error (Boroojeny et al., 2024; Sedghi et al., 2018; Senderovich et al., 2022). We are also inspired by seminal works that revealed the spectral properties of the Koopman operator (Rowley et al., 2009; Mezić, 2013; Mezic, 2020). Indeed, if one clips the spectrum of a learned Koopman matrix, this simple approach can help enforce stability when learning nonlinear systems.
We designed and carried out thorough experiments involving diverse applications and benchmark datasets to investigate the effectiveness of SC and compared its performance against strong baselines. We specifically compared SC with three baselines: CG (Boots et al., 2007), WLS (Huang et al., 2016), and SOC (Mamakoukas et al., 2020). We first evaluated SC’s ability to learn stable autonomous linear systems by considering the challenge of learning dynamic textures from videos from two benchmark datasets (Chan and Vasconcelos, 2005; Hadji and Wildes, 2018). We also evaluated SC’s ability to learn systems with control inputs by evaluating its performance when learning the dynamic model of a Franka Emika Panda robotic arm from a standardized dataset (Gaz et al., 2019). Our results across all experiments demonstrate that SC achieves the lowest reconstruction error and generates stable long-term predictions. Most notably, SC is able to achieve this level of performance while being orders-of-magnitude faster in learning provably-stable linear systems.
@@
We also investigated if SC can be combined with Koopman operators to learn complex nonlinear systems. Specifically, we evaluated SC’s ability to learn stable linear systems in the lifted space that encode complex dexterous manipulation skills for a multi-fingered robotic hand (Han et al., 2023). Our analyses reveals that unconstrained learning produces unstable linear systems if the training data includes unsuccessful demonstrations or when the task horizon is truncated. We are motivated by the fact that such characteristics are commonly observed when robot learning data is collected from humans (Qin et al., 2022; Shaw et al., 2023; Bahl et al., 2022). Our results show that SC can be readily combined with Koopman operators to learn stable nonlinear systems in mere seconds, while remaining robust to the undesirable characteristics in the training data. We also find that unstable and erratic robot behavior can arise due to a subset of the underlying dynamic modes (see Schmid (2022)). We find that clipping the spectrum of the learned Koopman matrix effectively stabilizes its rollouts and enables the robot to safely complete the task. These findings are in line with the well-established connections between Koopman modes and the state-space geometry of the underlying dynamical systems (Rowley et al., 2009; Mezić, 2013; Mezic, 2020).
@@ 2 Problem Formulation
Consider a discrete-time autonomous linear system as
$\mathrm{x}_{t+1}=A\mathrm{x}_{t},$ (1)
where $\mathrm{x}_{t}\in\mathcal{X}\subset\mathbb{R}^{n}$ is the system state at time $t$, and $A\in\mathcal{S}_{A}\subset\mathbb{R}^{n\times n}$ is the system evolution matrix. Here, we use $\mathcal{S}_{A}$ to denote the solution space of the matrices $A$.
Further, suppose we have a dataset $D=[\mathrm{x}_{1},\mathrm{x}_{2},\cdots,\mathrm{x}_{T}]$. Let $Y=[\mathrm{x}_{2},\mathrm{x}_{3},\cdots,\mathrm{x}_{T}]\in\mathbb{R}^{n\times T-1},X=[\mathrm{x}_{1},\mathrm{x}_{2},\cdots,\mathrm{x}_{T-1}]\in\mathbb{R}^{n\times T-1}$, then to learn the linear system from the dataset $D$ is to minimize the regression objective as follow:
$A=\arg\min_{\mathcal{S}_{A}}||Y-AX||_{F}^{2},$ (2)
Where $||\cdot||_{F}$ is the matrix Frobenius norm. Now, optimizing the $A$ matrices amounts to solving a least-square regression problem (Montgomery et al., 2021; Seber and Lee, 2012).
However, optimizing the Equation. 2 does not impose any constraints on the matrix $A$, such as the system stability that is crucial for the long-term prediction. Long-term prediction is particularly important for various applications, such as ensuring safety in robotic manipulation (Han et al., 2023, 2024). Therefore, we need to consider the stability guarantee constraint (Boots et al., 2007; Huang et al., 2016). Let $\{\lambda_{i}(A)\}^{n}_{i=1}$ be the eigenvalues of the matrix $A$, such matrix $A$ is stable if it has $\max\{\lambda_{i}(A)\}^{n}_{i=1}<=1$ (i.e., all the eigenvalues have a magnitude of one at most). Now, we introduce $\mathbb{S}_{A}=\{A\in\mathcal{S}_{A}|\max\{\lambda_{i}(A)\}^{n}_{i=1}<=1\}$ to denote the solution space where matrix $A$ is stable.
Learning Stable Linear Systems from Data:
Our goal is to learn the stable linear systems from the dataset $D$. Therefore, the objective in Equation. 2 turns into
$A=\arg\min_{\mathbb{S}_{A}}||Y-AX||_{F}^{2},$ (3)
which is highly non-convex (Boyd and Vandenberghe, 2004; Mamakoukas et al., 2020). In practice, obtaining an analytical solution to Equation 3 is impractical and necessitates approximate methods.
3 Spectrum Clipping for Stability
In this paper, we evaluate a simple method, which we refer to as Spectrum Clipping (SC), that effectively learns stable linear systems while being significantly more computational efficient.
@@ 3 Spectrum Clipping for Stability
In this method, we first compute the system evolution matrix $A_{ls}$ by optimizing the least square regression (Eqn. 2). Then, we perform eigen decomposition of the matrix $A_{ls}$
$A_{ls}=M\times\Lambda\times M^{-1}$ (4)
where $\Lambda$ is a diagonal matrix with all the eigenvalues of $A_{ls}$ on its diagonal, and $M$ is the matrix of corresponding eigenvectors (modal matrix).
There might be some defective eigenvalues whose algebraic multiplicity is larger than its geometric multiplicity. In this case, we may assign linear dependent eigenvectors to them but still keep the $\Lambda$ matrix diagonal. Although this may result in a non-invertible modal matrix as well as a slight numerical difference between the original matrix $A_{ls}$ and the reconstructed matrix $\bar{A_{ls}}=M\times\Lambda\times M^{-1}$, the deviation is negligible w.r.t the task performance in our experiments (an exception is discussed in section. 5).
For those unstable eigenvalues (i.e., eigenvalues whose norm is larger than 1) in matrix $\Lambda$, we clip their values in one of the following two ways.
Spectrum Clipping to 1
we clip them to exactly $1$ and constitute a new matrix $\bar{\Lambda}$. Finally, we multiply $\bar{\Lambda}$ with the original eigenvector matrices and get a modified matrix $\bar{A_{sc}}=M\times\bar{\Lambda}\times M^{-1},\bar{A_{sc}}\in\mathbb{S}_{A}$. This strategy allows us to obtain a linearly stable matrix that maintains maximum expressivity, which is also beneficial for long-term prediction. For instance, when learning dynamic textures, expressiveness is essential for accurate prediction beyond the training data horizon.
Spectrum Clipping to 1 - $\bm{\varepsilon}$, $\bm{\varepsilon\in}$ (0, 1]
By selecting $\varepsilon>0$, e.g., $\varepsilon=10^{-5}$, the resulting system stability meets the criteria of well-known Lyapunov stability [^1] All eigenvalues have magnitudes strictly less than one.. However, this choice might limit the system’s expressivity, and detailed discussions are included in the Appendix. D. Similarly, we then recompose the clipped $\bar{\Lambda}$ matrix with modal matrix $M$ and obtain $\bar{A_{sc}}$.
@@ 4 Empirical Analyses
In this section, we evaluate SC in terms of computation efficiency, stability, and expressivity when learning LDS from data and compare its performance against existing methods. Note that for all experiments, we clip the spectrum to one (see Appendix. D for the effects of different $\varepsilon$ values).
@@ 4.1 Experimental Design
We conduct all experiments on Alienware Aurora R13, with 10-core 12th Gen Intel Core i7-12700KF 1.0-GHz CPU and a 32G RAM.
@@
We compare the performance of SC against four baselines: i) LS: vanilla least squares, ii) (CG): constraint generation (Boots et al., 2007),iii) WLS: weighted least squares (Huang et al., 2016), and iv) SOC: a characterization approach (Mamakoukas et al., 2020, 2023).
@@
USCD (Chan and Vasconcelos, 2005): A dataset of 254 highway traffic videos. Each video has 48-52 frames and the size of each frame is $48\times 48$. To lower the dimension, we apply the SVD method on each sequence to obtain a subspace dimensions $r\in\{3,30\}$, as done by Mamakoukas et al. (2020).
DTDB (Hadji and Wildes, 2018): A dataset of 285 dynamic texture movements, such as turbulence, wavy motions, etc. Each video consists of varying frames, and we also apply the SVD method to lower the state dimension to $r=300$.
Note that due to the high dimension of the state space in the DTDB datasets, only LS, SOC, and SC are applicable. Both CG and WLS involve solving a Kronecker product, which has a space complexity of $O(d^{4})$ ($d$ is the state space dimension), leading to memory overflow.

### v4

@@ Abstract
When learning stable linear dynamical systems from data, three important properties are desirable: i) predictive accuracy, ii) provable stability, and iii) computational efficiency. Unconstrained minimization of reconstruction errors leads to high accuracy and efficiency but cannot guarantee stability. Existing methods to enforce stability often preserve accuracy, but do so only at the cost of increased computation. In this work, we investigate if a straightforward approach can simultaneously offer all three desiderata of learning stable linear systems. Specifically, we consider a post-hoc approach that manipulates the spectrum of the learned system matrix that was computed using unconstrained least squares. We call this approach spectrum clipping (SC) as it involves eigen decomposition and subsequent reconstruction of the system matrix after clipping any eigenvalues that are larger than one to one (without altering the eigenvectors). Through comprehensive experiments involving two different applications and publicly available benchmark datasets, we show that this simple technique can efficiently learn highly-accurate linear systems that are provably-stable. Notably, we find that SC can match or outperform strong baselines while being orders-of-magnitude faster. We also show that SC can be readily combined with Koopman operators to learn stable nonlinear dynamics, such as those underlying complex dexterous manipulation skills involving multi-fingered robotic hands. Finally, we find that SC can learn stable robot policies even when the training data includes unsuccessful or truncated demonstrations.
Keywords: Linear Dynamical Systems, Stability, System Identification
@@ 1 Introduction
In spite of the growing dominance of deep neural networks, modeling and learning of linear dynamical systems (LDS) remain relevant due to the availability of analytical solutions, significantly lower computational burden, and ease of analyses and inspection. LDS continue to find use in diverse applications such as computer vision [1, 2, 3, 4] and time series prediction [5, 6]. Even when the underlying dynamics are nonlinear, researchers are increasingly utilizing tools from linear systems theory by leveraging the Koopman Operator theory [7, 8, 9] to approximately represent nonlinear systems as linear systems in a higher-dimensional lifted state space. These techniques have further increased the relevance and importance of learning LDS in complex applications, such as legged robotics [10, 11], robotic manipulation [12, 13, 14, 15], atmospheric sciences [16, 17, 18], and fluid dynamics [19, 20].
@@
When time-series data associated with system evolution is available, learning LDS can be considered as a form of self-supervised learning and can be solved efficiently by minimizing a regression objective [21]. However, such unconstrained optimization methods tend to ignore important properties of the underlying systems, such as stability. Consequently, even if the underlying system is known to be stable, the stability of learned system could be highly sensitive to the training data distribution, potentially leading to unstable solutions and catastrophic long-term predictions.
@@
A number of methods have been developed to enforce stability of learned LDS and reduce sensitivity to training data [22, 23, 2, 24, 25, 26, 27]. These methods take one of three approaches: a) minimizing the regression objective under stability constraints [22, 23], b) iterating between solving an unconstrained problem and refining the solution under stability constraints [2, 24], and c) iterating between characterizing the stable matrix and optimizing the characterizations with the regression objective [25, 26, 27]. Regardless of the underlying approach, these methods inevitably introduce a considerable computational burden, potentially countering one of the primary advantages of leveraging linear techniques.
@@
In this work, we investigate a simple yet surprisingly effective method, Spectrum Clipping (SC), to learn stable LDS from data while introducing negligible additional computations. SC involves a straightforward implementation: 1) compute the system matrix via unconstrained least-squares, 2) perform eigen decomposition of the system matrix and extract all eigenvalues, 3) clip eigenvalues with magnitudes larger than one to one, and 4) reconstruct the system evolution matrix using the clipped eigenvalues and the unchanged eigenvectors. Our investigation is inspired by prior works that have revealed the surprising effectiveness of linear techniques: controlling the singular values of linear layers improves training stability and generalization error [28, 29, 30]; retraining only the last linear layer improves robustness to spurious correlations [31]; and Low-Rank Adaptation (LoRA) dramatically improves efficiency [32].
SC enforces stability by construction without requiring constrained optimization or iterative updates. In contrast to existing approaches discussed above, SC can efficiently scale to high-dimensional systems (e.g., image prediction). Further, seminal works on the Koopman operator [33, 19, 9] reveal fundamental connections between the spectrum of the Koopman matrix and desirable properties in learned nonlinear systems. SC can be trivially combined with Koopman operators to enhance the predictive accuracy and stability of learned nonlinear dynamical systems (e.g., robotic systems).
We conducted comprehensive experiments involving benchmark datasets across different domains (two video prediction datasets [1, 34] and one robot model learning dataset [35]) and strong baselines (CG [2], WLS [24], and SOC [26]). Our results reveal that SC consistently achieves the lowest reconstruction error, and generates stable and more accurate long-term predictions. Notably, SC is able to achieve this level of performance while being orders-of-magnitude faster.
@@
Additional experiments show that SC can be readily combined with Koopman operators to learn stable nonlinear systems (such as those that underlie complex dexterous manipulation skills for a multi-fingered robotic hand [13]) in mere seconds. Importantly, we show that SC enables the learned system to remain robustly stable and effective even the training data includes unsuccessful demonstrations or when the task horizon is truncated (characteristics are commonly observed when robot learning data is collected from humans [36, 37, 38]). In contrast, we find that unconstrained learning produces unstable LDS that lead to erratic and ineffective robot behaviors when such challenging characteristics are present in the training data.
In summary, our core contributions include: i) a simple post-hoc approach to learn stable linear dynamical systems, ii) evidence for its surprising ability to simultaneously offer reliable long-term predictions, high computational efficiency, and stability guarantees, iii) the ability to trade off between expressivity and stability, iv) an extension leveraging Koopman operators to accommodate stable nonlinear systems, and v) robustness against unsuccessful and truncated demonstrations.
@@ 2 Problem Formulation
Consider a stable discrete-time autonomous linear dynamical system $\mathrm{x}_{t+1}=A\mathrm{x}_{t}$, where $\mathrm{x}_{t}\in\mathcal{X}\subset\mathbb{R}^{n}$ is the system state at time $t$, and $A\in\mathcal{S}_{A}$ is the unknown system matrix, and $\mathcal{S}_{A}\subset\mathbb{R}^{n\times n}$ is the subspace that contains stable matrices: $\mathcal{S}_{A}\triangleq\{A\in\mathbb{R}^{n\times n}|\max\{|\lambda_{i}(A)|\}^{n}_{i=1}<=1\}$, where $\{\lambda_{i}(A)\}^{n}_{i=1}$ are the eigenvalues of the matrix $A$, and $|\lambda|$ is the $|\lambda|\coloneqq\sqrt{\text{Re}(\lambda)^{2}+\text{Im}(\lambda)^{2}}$.
Unconstrained Learning: Suppose we have a dataset $D=[\mathrm{x}_{1},\mathrm{x}_{2},\cdots,\mathrm{x}_{T}]$ containing the time evolutions of the above stable LDS. Defining $Y=[\mathrm{x}_{2},\mathrm{x}_{3},\cdots,\mathrm{x}_{T}]\in\mathbb{R}^{n\times T-1},X=[\mathrm{x}_{1},\mathrm{x}_{2},\cdots,\mathrm{x}_{T-1}]\in\mathbb{R}^{n\times T-1}$, allows one to express the problem of learning the above LDS from a dataset $D$ as the following solving standard least-square regression problem [39, 40]:
$\hat{A}_{LS}=\arg\min_{\Theta}||Y-\Theta X||_{F}^{2},$ (1)
where $\Theta\in\mathbb{R}^{n\times n}$ are the learnable parameters and $||\cdot||_{F}$ is the matrix Frobenius norm. However, optimizing the Eq. (1) ignores the constraints on the matrix $A$, such as the system stability that is crucial for the accurate long-term prediction.
Learning Stable LDS: Our goal is to ensure that the learned system respects the stability constraints of the underlying unknown LDS, despite learning from the limited dataset $D$. To this end, we can rewrite the optimization problem Eq. (1) to include stability constraints:
$\hat{A}_{S}=\arg\min_{\Theta\in\mathcal{S}_{A}}||Y-\Theta X||_{F}^{2},$ (2)
While solving the optimization problem in Eq. (2) can yield the constrained optimal solution which is guaranteed to be stable, it is known to be highly non-convex and requires approximations [41, 26].
3 Spectrum Clipping: Enforcing Stability of Learned LDS
We evaluate a simple method, which we refer to as Spectrum Clipping (SC), to ensure the stability of learned LDS without having to directly solve the constrained optimization in Eq. (2).
@@ 3 Spectrum Clipping: Enforcing Stability of Learned LDS
Eigen Decomposition of Unconstrained Solution: First, we compute the system evolution matrix $\hat{A}_{LS}$ by optimizing the least square regression (Eq. (1)). Second, we perform eigen decomposition of the least squares solution: $\hat{A}_{LS}=M\Lambda M^{-1}$, where $\Lambda$ is a diagonal matrix with all the eigenvalues of $\hat{A}_{LS}$ on its diagonal, and $M$ is the modal matrix with the corresponding eigenvectors.
Spectral Clipping: Next, we identify all unstable eigenvalues in $\Lambda$ and clip their magnitudes to $1$: $\bar{\Lambda}_{i}=1,\forall i=j,\text{such that}|\Lambda_{j}|>1$ , and leave all other eigenvalues in $\Lambda$ unaltered: $\bar{\Lambda}_{i}=\Lambda_{i},\forall i=j,\text{such that}|\Lambda_{j}|\leq 1$. Finally, we reconstruct the system matrix based on the new $\bar{\Lambda}$ and the unaltered modal matrix $M$ to produce SC’s estimate of the system matrix: $\hat{A}_{SC}\triangleq M\bar{\Lambda}M^{-1}$.
Note that the above post-hoc clipping procedure enforces stability constraints by ensuring that $\hat{A}_{SC}\in\mathcal{S}_{A}$ without requiring constrained optimization. Indeed, it is natural to suspect that this seemingly-naive approach would result in significantly poorer, albeit stable, approximations and worse prediction accuracies. However, as our experiments demonstrate, SC surprisingly leads to similar, if not better, accuracy compared to existing constrained optimization techniques.
Ensuring Lyapunov Stability: Enforcing Since SC clips eigenvalues to 1, the resulting system may not be strictly stable in the sense of Lyapunov. If desired, SC can be trivially extended to ensure Lyapunov stability by merely clipping the magnitude of the eigenvalues to marginally less than $1$. We formalize this by clipping the magnitudes to $1-\varepsilon$ (instead of $1$), where $\varepsilon\in[0,1]$ and Lyapunov stability is ensured when $\varepsilon>0$.
Expressivity-Stability Tradeoff: We note that the choice of $\varepsilon$ above represents an important tradeoff between the system’s expressivity and stability. In particular, choosing $\varepsilon=0$ will yield marginally-stable but highly-expressive systems. In contrast, choosing $\varepsilon>0$ will ensure strictly-stable system, with increasing values of $\varepsilon$ yielding faster exponential convergence and diminishing expressivity (see Appendix. E for experimental validation).
Dealing with Non-diagonalizable Unconstrained Solutions: A potential concern about the above approach is that $\hat{A}_{LS}$ could be non-diagonalizable. However, it is well known that one can always first find a diagonalizable matrix arbitrarily close to the original non-diagonalizable matrix (see Lemma. 3.1).
Lemma 3.1.
For any $\gamma>0$ and any non-diagonalizable matrix $\hat{A}_{LS}$, there exists a diagonalizable matrix $\tilde{A}_{LS}\in\mathbb{R}^{n\times n}$ s.t. $||\hat{A}_{LS}-\tilde{A}_{LS}||_{1}<\gamma$.
Proof.
Lemma 3.1 can be proved by the fact that the set of real matrices that are diagonalizable over $\mathbb{C}$ is dense in $\mathbb{R}^{n\times n}$. The details can be found from the matrix perturbation theory [42, 43]. ∎
The above result can be trivially extended to show that the evolution of the system with a non-diagonalizable matrix and that of the system with the arbitrarily-close diagonalizable matrix can be made arbitrarily close (see proof in Appendix. A). As such, the SC method can be applied to the arbitrarily-close diagonalizable matrix $\tilde{A}_{LS}$ to enforce stability of systems with non-diagonalizable matrices without significant approximation errors.
@@ 4 Empirical Analyses
In this section, we evaluate SC in terms of computation efficiency, stability, and expressivity when learning LDS from data and compare its performance against existing methods. Note that for all experiments, we clip the spectrum to $1$ (see Appendix. E for the effects of different $\varepsilon$ values).
@@ 4.1 Experimental Design
We conducted all experiments on Alienware Aurora R13, with 10-core 12th Gen Intel Core i7-12700KF 1.0-GHz CPU and a 32G RAM.
@@
i) LS: vanilla least squares, ii) (CG): constraint generation [2], iii) WLS: weighted least squares [24], and iv) SOC: a characterization approach [26, 27].
@@
UCSD [1]: A dataset of 254 highway traffic videos. Each video has 48-52 frames and the size of each frame is $48\times 48$. To lower the dimension, we apply the SVD method on each sequence to obtain a subspace dimensions $r\in\{3,30\}$, as similarly done in [26].
DTDB [34]: A dataset of 285 dynamic texture movements, such as turbulence, wavy motions, etc. Each video consists of varying frames, and we also apply the SVD method to lower the state dimension to $r=300$.